if a+b+c = 2 theta, then cos theta + cos (theta – a) + cos (theta – b) + cos (theta – c) is

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Let a,b,c be 60

a+b+c= 2*90 =180

So required solution is 3*(cos 30)

Regards